8 c) Draw the Ferrar graph of the following partitions:

i) 20 = 9+6+3+2

ii) 28 = 10+9+8+1

Also, write down the conjugate partitions of each of these partitions.


1
Expert's answer
2021-04-23T13:11:58-0400

(i) 20=9+6+3+2




The conjugate paratition is =9+3+6+2


(ii) 28=10+9+8+1




The conjugate paratition is=10+8+9+1